OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form

OpenStudy (aroub):

y=mx+b (-4,2) 2=-(-4)+b 2=4+b -4+2=b -2=b y=-x-2

OpenStudy (anonymous):

since the line is parallel, the slope would be the same i.e. '-1'. so the eqn would be : (y-2)/(x-(-4))=-1